# Revisiting institutional punishment in the $N$-person prisoner’s dilemma

### v2

@@ Abstract
The conflict between individual and collective interests makes fostering cooperation in human societies a challenging task, requiring drastic measures such as the establishment of sanctioning institutions. These institutions are costly because they have to be maintained regardless of the presence or absence of offenders. Here, we propose realistic improvements to the standard $N$-person prisoner’s dilemma formulation with institutional punishment by eliminating overpunishment, requiring a minimum number of contributors to establish the sanctioning institution, and sharing the cost among them once this minimum number is reached. In addition, we focus on large groups or communities for which sanctioning institutions are ubiquitous. Using the replicator equation framework for an infinite population, we find that by sufficiently fining players who fail to contribute either to the public good or to the sanctioning institution, a population of contributors immune to invasion by these free riders can be established, provided that the contributors are sufficiently numerous. In a finite population, we use finite-size scaling to show that, for some parameter settings, demographic noise helps to fixate the strategy that contributes to the public good but not to the sanctioning institution even for infinitely large populations when, somewhat counterintuitively, its proportion in the initial population vanishes with a small power of the population size.
I Introduction
@@
Understanding and promoting cooperation in human societies has been declared one of the greatest challenges for science in the twenty-first century Kennedy_2005 . The main issue is the conflict between individual and collective interests: either individuals contribute to public goods at a personal cost, or they free ride on the efforts of others Hardin_1968 ; Axelrod_1984 . This conflict between different levels of selection, which generally arises when considering social traits, was already noted by Darwin in his discussion of courage and self-sacrifice in humans, since these are traits that are counter-selected within a social group, but increase the group’s chances of survival in the case of inter-group fighting Darwin_1871 . This type of social dilemma has been extensively studied within the framework of evolutionary game theory Maynard_1982 .
@@
The usual evolutionary game theoretical framework for studying cooperation is the $N$-person prisoner’s dilemma Fox_1978 . In a play group of $N>1$ individuals, each individual can decide whether to contribute a fixed amount $c>0$ to the common pool (public goods). This amount is multiplied by a factor of $r>1$ and then divided among the $N-1$ other players. Following the terminology of the $2$-person prisoner’s dilemma, we refer to individuals that contribute to the common pool as cooperators and those that do not as defectors. Clearly, in this scenario cooperators are doomed to extinction unless there is some positive assortment between them Hamilton_1975 , which is best illustrated by the green-beard effect Hamilton_1964 , or repeated interactions that allow the emergence of reputation and reciprocity Wang_2023 ; Lai_2024 .
@@
An obvious way to promote cooperation would be to punish defectors. In fact, this works well when it comes to eliminating defectors, but since punishment is costly, peer punishers (who also contribute to public goods) are also eliminated, leaving the resulting population of cooperators vulnerable to invasion by defectors Boyd_1992 . In this scenario, a peer punisher imposes a fixed penalty on each defector. Introducing second-order punishment, i.e., punishing those who do not punish, leads to the elimination of both defectors and cooperators. However, since second-order punishment is ineffective in the absence of defectors, cooperators may take over the population due to random drift, leaving it vulnerable to invasion by defectors again. One way out of this conundrum is to assume that the game is noncoercive, i.e., some players may abstain from participating, and to set the payoff of nonparticipants to an ad hoc value greater than the payoff received by a defector in a group of defectors, but smaller than the payoff received by a pair of cooperators Fowler_2005 ; Hauert_2007 . This guarantees that a population of defectors can be invaded by nonparticipants, which in turn can be invaded by cooperators, which can be invaded by defectors again, creating a sort of rock-paper-scissors game scenario Hannelore_2010 .
@@
However, at least from the perspective of (modern) human social organization, a more natural way to deal with defectors of all kinds is to introduce pool or institutional punishment, where each punisher contributes a fixed amount in advance to a sanctioning institution Sigmund_2010 ; Szolnoki_2011 ; Perc_2012 . In particular, institutional punishers not only contribute $c$ to the public goods, but also advance an amount $\gamma$ to a sanctioning institution. Defectors are fined a fixed amount $\beta$ and cooperators are fined a smaller amount $\alpha\beta$ with $\alpha\in[0,1]$ for not contributing to the the sanctioning institution. This mimics the way contemporary societies are organized Ostrom_1990 . In particular, not participating in the game (e.g., not recognizing the right of the sanctioning institution to punish) is not an option.
@@
Given the undeniably important role that institutional punishment plays in curbing defection in the real world, we consider here some realistic improvements to the standard $N$-person prisoner’s dilemma formulation with pool punishment Sigmund_2010 . In particular, we eliminate overpunishment (i.e., penalties for different types of offenses are fixed and do not depend on the wealth of the sanctioning institution Dercole_2013 ), but account for the high costs of creating and maintaining a sanctioning institution by requiring a minimum number of punishers to establish it. However, the maintenance costs are divided among the punishers once this minimum number is reached. The sanctioning institution must be maintained regardless of the presence or absence of offenders Perc_2012 . In addition, and in contrast to most studies of public goods games that consider small groups of players, we focus on large groups (i.e., communities) for which sanctioning institutions are pervasive.
@@
In the infinite population limit, where the game can be studied using the replicator equation framework Hofbauer_1998 , we find that the all-defectors solution is always stable and the all-cooperators solution is always unstable. An equilibrium solution where the population consists only of institutional punishers is stable if the per capita cost to punishers of maintaining the sanctioning institution (the maximum per capita cost is $\gamma$) is less than the fine to cooperators ($\alpha\beta$) and this cost plus the cost $c$ of contributing to public goods is less than the fine to defectors ($\beta$). In this bistability scenario, there is a lower value of the frequency of punishers in the population at large that guarantees the maintenance of the sanctioning institution and, consequently, the disappearance of both types of offenders (i.e., cooperators and defectors). This threshold has a non-trivial dependence on the size of the community. These findings are consistent with the results of other implementations of institutional punishment where both cooperators and defectors are punished Szolnoki_2011 ; Perc_2012 ; Dercole_2013 .
For finite populations, the game is studied using the imitation stochastic dynamics, which reproduces the results of the replicator equation framework in the infinite population limit Fontanari_2024b ; Traulsen_2005 ; Sandholm_2010 . In the case the fine to cooperators is too small, so they win over the punishers in the deterministic regime, we use finite-size scaling Binder_1985 ; Privman_1990 ; Campos_1999 to show that demographic noise helps to fixate cooperators even for infinitely large populations, provided, somewhat counterintuitively, that the initial fraction of cooperators vanishes with some small power of the population size.
@@
The remaining sections are organized as follows. In Section II, we present a variant of the $N$-person prisoner’s dilemma that more realistically models both the maintenance and the operation of sanctioning institutions. In particular, in this section we present the instantaneous payoffs of the three strategies that players can adopt, namely cooperators, defectors, and institutional punishers, and describe the stochastic imitation dynamics that govern the evolution of strategy frequencies for finite populations. In Section III we study the game for infinitely large populations using the replicator equations framework: we find the equilibrium solutions and the conditions for local stability, and explore the dynamics by numerically solving the replicator equations and representing the orbits in the simplex. In Section IV we study the finite population version of the game, using Monte Carlo simulations to implement the stochastic imitation dynamics and finite-size scaling techniques to infer fixation probabilities and mean time to fixation in the limit of very large populations. In Section V, we recapitulate our main findings and present some closing observations.
II The model
@@
The focal individual’s payoff depends on her own strategy as well as the strategies of the other members of her group. Let us consider each possibility separately.
@@
    $F_{C}(I+1,J,K)$ $=$ $\frac{(I+K)rc}{N-1}-c$ (1)
    $-\alpha\beta\Theta(K-K_{m})$
    where the Heaviside function is $\Theta(x)=1$ if $x\geq 0$ and 00 otherwise. Here we consider the so-called others-only scenario, in which the amplified contribution $rc$ is shared by all other participants in the play group, but not by the contributor Hannelore_2010 (see also Hamilton_1975 ). This is the reason why the numerator of the first term on the rhs of Eq. (1) is missing one contributor (the total number of contributors is $I+1+K$). Moreover, each amplified contribution is shared only by $N-1$ members of the group, since the contributors do not benefit from their own contributions. Hence the factor $N-1$ in the denominator of this term. The second term on the rhs of Eq. (1) is the focal individual’s contribution to public goods, and the third term is the penalty for not contributing to the sanctioning institution. This penalty is a fraction $\alpha\in[0,1]$ of the penalty applied to defectors (see below) and it is only effective if there are at least $K_{m}<N$ individuals (i.e., punishers) contributing to the sanctioning institution. Once this minimum number of punishers is reached, the penalty term becomes constant, in contrast to previous formulations of institutional punishment where the penalty increases linearly with the number of individuals contributing to the sanctioning institution Hannelore_2010 ; Sigmund_2010 , resulting in overpunishment Dercole_2013 . In fact, the constant punishment term with $K_{m}=1$ was proposed in the original study that showed the non-necessity of overpunishment for the establishment of public goods contributors in the institutional punishment scenario Dercole_2013 .
@@
    $F_{D}(I,J+1,K)$ $=$ $\frac{(I+K)rc}{N-1}$ (2)
    $-\beta\Theta(K-K_{m}).$
@@
    $F_{P}(I,J,K+1)$ $=$ $\frac{(I+K)rc}{N-1}-c-\gamma$ (3)
    $+\gamma\Theta(K+1-K_{m})$
    $-\gamma\frac{K_{m}}{K+1}\Theta(K+1-K_{m}).$
    Only the last term in the rhs of this equation needs explanation. There is a minimum cost of creating the sanctioning institution, which can only be covered by the contribution of $K_{m}$ punishers. This is similar to the variant of the NPD where a minimum number of cooperators is needed to produce the public goods Pacheco_2009 . Once this threshold cost is reached, the punishers optimize their payoffs by sharing among them the costs of maintaining the sanctioning institution, a feature that makes this variant of the NPD similar to the $N$-person snowdrift game Zheng_2007 ; Santos_2012 ; Fontanari_2024a . Besides the introduction of the threshold $K_{m}$, another important difference between our approach and that of Ref. Dercole_2013 is that punishers always contribute to the sanctioning institution, i.e., they contribute even if there are no defectors and cooperators. In fact, their contribution does not depend on the number of offenders at all. These assumptions are more consistent with the way sanctioning institutions are maintained by contributors in the real world.
@@
which is the difference between the payoff obtained by a single defector in a group of cooperators (i.e., $rc$) and the payoff obtained by a single punisher in a group of defectors (i.e., $-c-\gamma$). The update (or not) of the focal individual concludes the time step $\delta t$ and the time variable $t$ is increased accordingly. The particular choice of switching probability given in Eq. (4) produces the replicator equation in the $M\to\infty$ limit Fontanari_2024b (see Traulsen_2005 ; Sandholm_2010 for other switching probabilities that produce the replicator equation in the large-population limit), provided that the time step is set to $\delta t=1/M$.
III The replicator equation formulation
Here we consider an infinitely large population (i.e., $M\to\infty$) consisting of a fraction $x$ of cooperators, a fraction $y$ of defectors and a fraction $z$ of punishers, with $x+y+z=1$. Only the expected payoffs go into the replicator equation Hofbauer_1998 , so we need to calculate them for each of the three strategies. In particular, the expected payoff of a cooperator is
$\pi^{C}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}F_{C}(I+1,J,K)\delta_{I+J+K,N-1}$ (6)
$=$ $rc(x+z)-c$
$-\alpha\beta\left[1-\sum_{K=0}^{K_{m}-1}\binom{N-1}{K}z^{K}(1-z)^{N-1-K}\right]$
@@ III The replicator equation formulation
$\pi^{D}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}F_{D}(I,J+1,K)\delta_{I+J+K,N-1}$ (7)
$=$ $rc(x+z)$
$-\beta\left[1-\sum_{K=0}^{K_{m}-1}\binom{N-1}{K}z^{K}(1-z)^{N-1-K}\right].$
@@
$\pi^{P}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}F_{P}(I,J,K+1)\delta_{I+J+K,N-1}$ (8)
$=$ $rc(x+z)-c$
$-\frac{\gamma K_{m}}{Nz}\left[1-\sum_{K=0}^{K_{m}-1}\binom{N}{K}z^{K}(1-z)^{N-K}\right]$
@@
Sanctioning institutions only make sense for very large play groups, otherwise peer punishment or small coalitions are more appropriate Ostrom_1990 . It is therefore of interest to obtain the expressions for the payoffs of the strategies in the limits $N\to\infty$ and $K_{m}\to\infty$, such that the ratio
@@
is nonzero. This can be easily achieved by replacing the binomial distributions by Gaussian distributions in Eqs. (6), (7) and (8) and doing the integrations over the number of punishers. The final result is
$\pi^{C}=\begin{cases}rc(x+z)-c&\text{if $z<\zeta$ }\\ rc(x+z)-c-\alpha\beta&\text{otherwise},\end{cases}$ (10)
$\pi^{D}=\begin{cases}rc(x+z)&\text{if $z<\zeta$ }\\ rc(x+z)-\beta&\text{otherwise},\end{cases}$ (11)
@@
$\pi^{P}=\begin{cases}rc(x+z)-c-\gamma&\text{if $z<\zeta$ }\\ rc(x+z)-c-\gamma\zeta/z&\text{otherwise}.\end{cases}$ (12)

### v3

@@ Abstract
The conflict between individual and collective interests makes fostering cooperation in human societies a challenging task, requiring drastic measures such as the establishment of sanctioning institutions. These institutions are costly because they have to be maintained regardless of the presence or absence of offenders. Here we revisit some improvements to the standard $N$-person prisoner’s dilemma formulation with institutional punishment in a well-mixed population, namely the elimination of overpunishment, the requirement of a minimum number of contributors to establish the sanctioning institution, and the sharing of its maintenance costs once this minimum number is reached. In addition, we focus on large groups or communities for which sanctioning institutions are ubiquitous. Using the replicator equation framework for an infinite population, we find that by sufficiently fining players who fail to contribute either to the public good or to the sanctioning institution, a population of contributors immune to invasion by these free riders can be established, provided that the contributors are sufficiently numerous. In a finite population, we use finite-size scaling to show that, for some parameter settings, demographic noise helps to fixate the strategy that contributes to the public good but not to the sanctioning institution even for infinitely large populations when, somewhat counterintuitively, its proportion in the initial population vanishes with a small power of the population size.
journal: Theory in Biosciences
1 Introduction
@@
Understanding and promoting cooperation in human societies has been declared one of the greatest challenges for science in the twenty-first century [1]. The main issue is the conflict between individual and collective interests: either individuals contribute to public goods at a personal cost, or they free ride on the efforts of others [2, 3]. This conflict between different levels of selection, which generally arises when considering social traits, was already noted by Darwin in his discussion of courage and self-sacrifice in humans, since these are traits that are counter-selected within a social group, but increase the group’s chances of survival in the case of inter-group fighting [4]. This type of social dilemma has been extensively studied within the framework of evolutionary game theory [5].
@@
The usual evolutionary game theoretical framework for studying cooperation is the $N$-person prisoner’s dilemma [6]. In a play group of $N>1$ individuals, each individual can decide whether to contribute a fixed amount $c>0$ to the common pool (public goods). This amount is multiplied by a factor of $r>1$ and then divided among the $N-1$ other players. Following the terminology of the $2$-person prisoner’s dilemma, we refer to individuals that contribute to the common pool as cooperators and those that do not as defectors. Clearly, in this scenario cooperators are doomed to extinction unless there is some positive assortment between them [7, 8], which is best illustrated by the green-beard effect [9], or repeated interactions that allow the emergence of reciprocity and reputation, which are not necessarily interrelated [10, 11] (see [12] for a recent review).
@@
An obvious way to promote cooperation would be to punish defectors [13, 14]. In fact, this works well when it comes to eliminating defectors, but since punishment is costly, peer punishers (who also contribute to public goods) are also eliminated, leaving the resulting population of (pure) cooperators vulnerable to invasion by defectors [13]. In the above scenario, a peer punisher imposes a fixed penalty on each defector. Introducing second-order peer-punishment, i.e., punishing cooperators for not punishing defectors, leads to the elimination of both defectors and cooperators. However, since second-order peer punishment is ineffective in the absence of defectors, because it is impossible to distinguish cooperators from punishers when there is no one to punish, cooperators may take over the population due to random drift, leaving it vulnerable to invasion by defectors again. One way out of this conundrum is to assume that the game is noncoercive, i.e., some players may abstain from participating, and to set the payoff of nonparticipants to an ad hoc value greater than the payoff received by a defector in a group of defectors, but smaller than the payoff received by a pair of cooperators [15, 16]. This guarantees that a population of defectors can be invaded by nonparticipants, which in turn can be invaded by cooperators, which can be invaded by defectors again, creating a sort of rock-paper-scissors game scenario [17]. We note, however, that the dramatic scenario we describe is only true for well-mixed populations [17, 18] and that considering spatial interactions such that players interact only with a fixed number of neighbors can lead to punishers becoming dominant. In fact, this limited set of interactions makes it possible to separate cooperators from punishers, who can then compete separately with defectors. As a result of their indirect territorial competition, cooperators are driven out, leaving only punishers [19, 20, 21, 22].
@@
At least from the perspective of (modern) human social organization, a more natural way to deal with defectors of all kinds is to introduce pool or institutional punishment, where each punisher contributes a fixed amount in advance to a sanctioning institution [18, 23]. The fact that such costs should be paid regardless of the presence or absence of offenders is what distinguishes institutional punishment from peer punishment, since peer punishers bear the cost of punishment only in the presence of defectors. In particular, institutional punishers not only contribute $c$ to the public goods, but also advance an amount $\gamma$ to a sanctioning institution. Defectors are fined a fixed amount $\beta$ and cooperators are fined a smaller amount $\alpha\beta$ with $\alpha\in[0,1]$ for not contributing to the sanctioning institution. Not participating in the game (e.g., not recognizing the right of the sanctioning institution to punish) is not an option so we do not include voluntary optional participation in our analysis. The main advantage of institutional punishment is that cooperators are easy to identify and punish even in the absence of defectors (they do not contribute to the sanctioning institution), which is not the case in the peer punishment scenario. This mimics the way contemporary societies are organized [24].
@@
Given the undeniably important role that institutional punishment plays in curbing defection in the real world, here we revisit some refinements to the the standard $N$-person prisoner’s dilemma formulation with pool punishment in a well-mixed population [18]. In particular, we eliminate overpunishment, i.e., penalties for different types of violations are fixed and do not depend on the wealth of the sanctioning institution, which is of course proportional to the number of punishers (see [25] for a similar approach in the context of peer punishment). Moreover, to account for the high cost of creating and maintaining a sanctioning institution we require a minimum number of punishers to establish it (see [26, 27] for a similar approach in the context of collective risk social dilemmas). In addition, to reduce the burden on punishers, we assume that once this minimum number is reached, the cost of maintenance will be shared among them (see [25, 28] for a similar approach in the context of peer punishment).
Here we consider three main research questions about institutional punishment in public goods games, the answers to which are our original contributions to this widely studied topic. First, we determine the effect of sharing the maintenance costs of the sanctioning institution on the stability of punishers. Second, in contrast to most studies of public goods games that consider small groups of players, we focus on large groups (i.e., communities) for which sanctioning institutions are ubiquitous. In analyzing the effect of group size $N$, we assume that the minimum number of punishers required to create the sanctioning institution increases linearly with $N$. Third, and more importantly, by simulating populations of finite size $M$, we study the effect of demographic noise on the predictions of the deterministic theory.
@@
In the infinite population limit, where the game can be studied using the replicator equation framework [29], we find that the all-defectors solution is always stable and the all-cooperators solution is always unstable. An equilibrium solution where the population consists only of institutional punishers is stable if the per capita cost to punishers of maintaining the sanctioning institution (the maximum per capita cost is $\gamma$) is less than the fine to cooperators ($\alpha\beta$) and this cost plus the cost $c$ of contributing to public goods is less than the fine to defectors ($\beta$). In this bistability scenario, there is a lower value of the frequency of punishers in the population at large that guarantees the maintenance of the sanctioning institution and, consequently, the disappearance of both types of offenders (i.e., cooperators and defectors). This threshold has a non-trivial dependence on the size of the community. In addition, we find that cost sharing increases the region of stability of the all-punishers equilibrium solution as well as the size of its domain of attraction.
For finite populations, the game is studied through simulations of the imitation stochastic dynamics, which reproduces the results of the replicator equation framework in the infinite population limit [30, 31, 32]. In the case the penalty $\alpha\beta$ to cooperators is too small, so they win over the punishers in the deterministic regime, we use finite-size scaling [33, 34, 35] to show that demographic noise helps to fixate cooperators even for infinitely large populations, provided, somewhat counterintuitively, that the initial fraction of cooperators vanishes with some small power of the population size $M$. Finite-size scaling is a concept from statistical physics that refers to the study of how the properties of a system change as a function of its finite size, especially near a threshold. The goal is to obtain the scaling forms and the exponents of the power-laws that govern the behavior of quantities of interest (e.g., the fixation probability and the mean fixation time) near a threshold, simulating large but finite system sizes. In this sense, it provides a systematic way to computationally study the effect of demographic noise in problems that do not lend themselves to analytical approaches such as Kimura’s diffusion approximation [36]. Finite-size scaling requires simulating the system for various sizes and rescaling the variables to obtain quantities that are independent of system size, resulting in the collapse of the data for the various system sizes into a single curve, the scaling function [34]. Besides studying the effect of demographic noise, we also use this technique to study the effect of group size $N$, since the deterministic theory gives closed analytical results for $N\to\infty$ but not for finite $N$.
@@
The remaining sections are organized as follows. In Section 2, we present a variant of the $N$-person prisoner’s dilemma that more realistically models both the maintenance and the operation of sanctioning institutions. In particular, in this section we present the instantaneous payoffs of the three strategies that players can adopt, namely cooperators, defectors, and institutional punishers, and describe the stochastic imitation dynamics that govern the evolution of strategy frequencies for finite populations. In Section 3 we study the game for infinitely large populations using the replicator equations framework: we find the equilibrium solutions and the conditions for local stability, and explore the dynamics by numerically solving the replicator equations and representing the orbits in the simplex. In Section 4 we study the finite population version of the game, using Monte Carlo simulations to implement the stochastic imitation dynamics and finite-size scaling techniques to infer fixation probabilities and mean time to fixation in the limit of very large populations. In Section 5, we recapitulate our main findings and present some closing observations.
2 The model
@@
The focal individual’s payoff depends on her own strategy as well as on the strategies of the other members of her group. Let us consider each possibility separately.
@@
    $F_{C}(I+1,J,K)=\frac{(I+K)rc}{N-1}-c-\alpha\beta\Theta(K-K_{m}),$ (1)
    where the Heaviside function is $\Theta(x)=1$ if $x\geq 0$ and 00 otherwise. Here we consider the so-called others-only scenario, in which the amplified contribution $rc$ is shared by all other participants in the play group, but not by the contributor [17] (see also [7]). This is the reason why the numerator of the first term on the rhs of Eq. (1) is missing one contributor (the total number of contributors is $I+1+K$). Moreover, each amplified contribution is shared only by $N-1$ members of the group, since the contributors do not benefit from their own contributions. Hence the factor $N-1$ in the denominator of this term. The second term on the rhs of Eq. (1) is the focal individual’s contribution to public goods, and the third term is the penalty for not contributing to the sanctioning institution. This penalty is a fraction $\alpha\in[0,1]$ of the penalty $\beta$ applied to defectors (see below) and it is only effective if there are at least $K_{m}<N$ individuals (i.e., punishers) contributing to the sanctioning institution. Once this minimum number of punishers is reached, the penalty term becomes constant, in contrast to previous formulations of institutional punishment where the penalty increases linearly with the number of individuals contributing to the sanctioning institution [17, 18], resulting in overpunishment. In fact, the constant punishment term with $K_{m}=1$ was proposed in the original study that showed the non-necessity of overpunishment for the establishment of public goods in the peer-punishment scenario [25].
@@
    $F_{D}(I,J+1,K)=\frac{(I+K)rc}{N-1}-\beta\Theta(K-K_{m}).$ (2)
@@
    $F_{P}(I,J,K+1)=\frac{(I+K)rc}{N-1}-c-\gamma\left[1-\left(1-\frac{K_{m}}{K+1}\right)\Theta(K+1-K_{m})\right].$ (3)
    Only the last term in the rhs of this equation needs explanation. There is a minimum cost of creating the sanctioning institution, which can only be covered by the contribution of $K_{m}$ punishers. This is similar to the variant of the NPD where a minimum number of cooperators is needed to produce the public goods [37], and is also a key feature of collective risk social dilemmas [26, 27]. Once this threshold cost is reached, the punishers optimize their payoffs by sharing among them the costs of maintaining the sanctioning institution, a feature that makes this variant of the NPD similar to the $N$-person snowdrift game [38, 39, 40], and which has also been considered in the context of peer punishment [25, 28]. It is interesting to mention a perhaps more realistic scenario where punishment is graduated or adaptive, growing with the incidence of defectors and with the harm resulting from failure to cooperate (see, e.g., [41, 42, 43]). Another interesting scenario considers that incentives (bonuses to cooperators) and punishments (fines to defectors) are given externally at no cost to the players, and the problem is to optimize the incentive and punishment schedules and intensities to achieve cooperation at the lowest cost [44, 45, 46].
@@
which is the difference between the payoff obtained by a single defector in a group of cooperators (i.e., $rc$) and the payoff obtained by a single punisher in a group of defectors (i.e., $-c-\gamma$). The update (or not) of the focal individual concludes the time step $\delta t$ and the time variable $t$ is increased accordingly. The particular choice of switching probability given in Eq. (4) produces the replicator equation in the $M\to\infty$ limit [30] (see [31, 32] for other switching probabilities that produce the replicator equation in the large-population limit), provided that the time step is set to $\delta t=1/M$.
3 The replicator equation formulation
Here we consider an infinitely large population (i.e., $M\to\infty$) consisting of a fraction $x$ of cooperators, a fraction $y$ of defectors and a fraction $z$ of punishers, with $x+y+z=1$. Only the expected payoffs go into the replicator equation [29], so we need to calculate them for each of the three strategies. In particular, the expected payoff of a cooperator is
$\pi^{C}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}x^{I}y^{J}z^{K}F_{C}(I+1,J,K)\delta_{I+J+K,N-1}$ (6)
$=$ $rc(x+z)-c-\alpha\beta\left[1-\sum_{K=0}^{K_{m}-1}\binom{N-1}{K}z^{K}(1-z)^{N-1-K}\right],$
@@ 3 The replicator equation formulation
$\pi^{D}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}x^{I}y^{J}z^{K}F_{D}(I,J+1,K)\delta_{I+J+K,N-1}$ (7)
$=$ $rc(x+z)-\beta\left[1-\sum_{K=0}^{K_{m}-1}\binom{N-1}{K}z^{K}(1-z)^{N-1-K}\right].$
@@
$\pi^{P}$ $=$ $\sum_{I,J,K=0}^{N-1}\frac{(N-1)!}{I!J!K!}x^{I}y^{J}z^{K}F_{P}(I,J,K+1)\delta_{I+J+K,N-1}$ (8)
$=$ $rc(x+z)-c-\frac{\gamma K_{m}}{Nz}\left[1-\sum_{K=0}^{K_{m}-1}\binom{N}{K}z^{K}(1-z)^{N-K}\right]$
@@
Sanctioning institutions only make sense for very large play groups, otherwise peer punishment or small coalitions are more appropriate [24]. It is therefore of interest to obtain the expressions for the payoffs of the strategies in the limits $N\to\infty$ and $K_{m}\to\infty$, such that the ratio
@@
is nonzero. This can be achieved by replacing the binomial distributions by Gaussian distributions in Eqs. (6), (7) and (8) and doing the integrations over the number of punishers (see A). The final result is
$\pi^{C}=\begin{cases}rc(x+z)-c&\text{if $z<\zeta$ }\\ rc(x+z)-c-\alpha\beta/2&\text{if $z=\zeta$ }\\ rc(x+z)-c-\alpha\beta&\text{if $z>\zeta$},\end{cases}$ (10)
$\pi^{D}=\begin{cases}rc(x+z)&\text{if $z<\zeta$ }\\ rc(x+z)-\beta/2&\text{if $z=\zeta$ }\\ rc(x+z)-\beta&\text{if $z>\zeta$},\end{cases}$ (11)
@@
$\pi^{P}=\begin{cases}rc(x+z)-c-\gamma&\text{if $z\leq\zeta$ }\\ rc(x+z)-c-\gamma\zeta/z&\text{if $z>\zeta$}.\end{cases}$ (12)